OpenStudy (anonymous):

could someone help me..what is 3/8 divided by 18

OpenStudy (anonymous):

0.0208333333

OpenStudy (anonymous):

i got 0.375???

OpenStudy (anonymous):

3 over 8 means 3 divided by 8 which equals 0.375. 0.375 divided by 18 then gives you 0.0208333333

OpenStudy (lgbasallote):

\(\Large \frac{\frac{3}{8}}{18} = (\frac{3}{8})(\frac{1}{18}) = (\frac{1}{8})(\frac{1}{6}) = \frac{1}{48}\)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Unless you wanted it in fractions like that ^ lol But in decimal form it is 0.0208333333

OpenStudy (anonymous):

thank you and yes i needed in fraction form..sorry but thank you